Check my work You can save $2,000 per year for the next four years in an account earning 7 percent per year. How much will you have at the end of the fourth year if you make the first deposit today? (Do not round intermediate calculations. Round your answer to 2 decimal places. (e.g., 32.16) Future value
Answer #1

FV = 2000*(1.07)4+2000*(1.07)3+2000*(1.07)2​+2000*(1.07)1

= 2621.6+2450+2290+2140 = 9501